Kendra has
$6.30
in nickels and quarters. If she has
12
more quarters than nickels, how many of each coin does she have?

Answer by solve_for_x(190)

Let x represent the number of quarters, and let y represent the number of nickels.

The total value of the coins is then:

0.25x + 0.05y = 6.30

Since she has 12 more quarters than nickels, you know that:

x = y + 12

Substituting for x in the first equation gives

0.25(y + 12) + 0.05y = 6.30

0.25y + 3 + 0.05y = 6.30

0.30y + 3 = 6.30

0.30y = 6.30 - 3

0.30y = 3.30

y = 3.30 / 0.30

y = 11

x = y + 12 = 11 + 12 = 23

Solution: She has 23 quarters and 11 nickels.

Check:

0.25(23) + 0.05(11) = 5.75 + 0.55 = 6.30